General Meeting Summary 7/29/21

From Cuawiki
Jump to: navigation, search

PARTICIPANTS: Aaron Brown, Arshak Asaturyan, Fernando Barbosa, Bogdan Wojtsekhowski, Brad Sawatzky, Brannon Semp, Carlos Munoz-Camacho, Carlos Yero, Charles Hyde, Rolf Ent, Hakob Voskanyan, Hamlet Mkrtchyan, Julie Roche, Marie Boer, Paulo Medeiros, Stephen Wood, Stephen Lassiter, Tanja Horn, Tristan Anderson, Vardan Tadevosyan, Vladimir Berdnikov, ALexandre Camsonne


HV DIVIDER UPDATES (Vladimir, Carlos Y., Fernando, Julie)

  • Linearity Tests vs. rate
  • No optical filter (glass, neutral density) with different laser (405nm) intensity
  • One optical filter and
  • Different optical filters and laser intensity FIXED
  • Results suggest that linearity of the divider is very good up to a few MHz
  • Currently measuring linearity of amplifier itself - inject signal
  • one optical filter (from UV to visible) for which one can set optical density to specify attentuation
  • Action items
  • Complete the linearity test of the amplifier
  • Work on next divider revision taking into account needs for new LV cabling strategy
  • Ship PMTs from OU to JLab - need to remove sockets; could also just ship the sockets


LV CABLING STRATEGY (Carlos M., IJCLab-Orsay team)

  • Have a working option for LV cabling that maintains easy access for changing PMTs - issue: many cables --> have been exploring alternatives to reduce number of cables
  • Follow up from last meeting: daisy-chaining (power PMTs in parallel) the LV vertically
  • 18 serial connectors
  • still need to check on voltage drop across chain to evaluate what is possible
  • Challenge with daisy-chaining: maintenance (access to PMTs and changing PMTs)
  • space available for board, connectors, etc.
  • connector plug/unplugging very difficult
  • Action items:
  • voltage drop across chain - from connection to connection (Bogdan)
  • explore different connector - Bogdan's suggestion of a flat/twisted connector - challenge here could be mechanical
  • Bogdan/Fernando provide information on connectors


DSG UPDATES (Aaron)

  • Hardware interlock system
  • Chiller interlock - includes a remote power controller (check if uses optical isolators - high failure rate in high power halls)
  • CAEN crate interlock - may want to add a label that CAEN CPU interlock remains in OPEN position
  • Action items:
  • check if remote power controller uses optical isolator
  • check on labeling for CAEN CPU


  • Action item (Fernando):
  • check on way to monitor/interlock anode current PMT or similar, e.g., DC level or so; may not need to do that for all 1200 channels

MECHANICAL (Paulo, Steven L.)

  • Much progress and most components are in procurement or in house
  • Remaining is cable routing and roof block modification to get cables into the SHMS hut
  • constraint: cables have already been fabricated
  • constraint on SHMS: with this configuration will be limited to ~39deg
  • NPS power supply requisition has been submitted


DDVCS WITH NPS (Marie)

  • Presentation on ideas for DDVCS with NPS


NEXT MEETING: THURSDAY AUGUST 26 AT 9AM ET